OpenStudy (anonymous):

Little sisters easy math questions. Convert the decimal to a percent. 0.007 A. 0.007% B. 0.07% C. 0.7% D. 7%

OpenStudy (anonymous):

To convert a decimal to a percent you multiply it by 100 So 0.007*100= 0.7%

OpenStudy (anonymous):

okay. what about a percent to a decimal?

OpenStudy (anonymous):

I think you would divide by 100 but I am not completely sure?

OpenStudy (anonymous):

C.

OpenStudy (alexandervonhumboldt2):

0.007*100=000.7%=0.7%

OpenStudy (alexandervonhumboldt2):

lol @robtobey gave a direct answer so funny

OpenStudy (anonymous):

yes but im asking a new question. how do you convert percent to a decimal? like this. Convert the percent to a decimal. 0.3% A. 0.003 B. 0.3 C. 3 D. 30

OpenStudy (anonymous):

or nah ig.

OpenStudy (anonymous):

the answer is D @EmoMuch?Yes.

OpenStudy (anonymous):

"%" can be viewed as "parts per hundred" 0.3/100 = 0.003

OpenStudy (anonymous):

From a percent to decimal you divide it by 100. 0.3/100= 0.003